# Which Factors Make Code LLMs More Vulnerable to Backdoor Attacks? A Systematic Study

### v1

@@ IV Results
We analyze backdoor attacks by varying data (RQ1), training (RQ2), and inference (RQ3) factors.
@@ IV-A1 Poisoning Rate
The poisoning rate measures the percentage of poisoned samples in the training dataset. We experiment with poisoning rates from 0.01% to 10% to cover previously overlooked low poisoning rates, evaluating three trigger types (fixed, grammar, and LLM-generated triggers) across three models (CodeT5, CodeT5+, and PLBART).
Starting from zero at 0.01% poisoning rate, the ASR surpasses 80% at just 0.1% poisoning rate for most experiments, except for the LLM-generated trigger on PLBART (9.4%). When the poisoning rate exceeds 0.5%, all experiments consistently achieve ASR above 95%. To better illustrate the growth trend of ASR, we increase sampling density between poisoning rates of 0.01% and 0.1%, as illustrated in Figure 4. For all experiments except the LLM-generated trigger on PLBART, the ASR increases from 0 to above 80% within an extremely narrow range - specifically between 0.03% and 0.09% (poisoning 6 more samples). Backdoor attacks remain a critical threat even below 0.1% poisoning rate—a range that is overlooked in existing works [9, 17, 4]. Researchers should expand backdoor evaluations to lower poisoning rates for a comprehensive understanding of their impact.
Figure 4 shows the FTR results under 0.01%-10% poisoning rates. Both CodeT5 and CodeT5+ maintain low FTR (below 0.2% and 0.1%) when poisoning rates are below 1%. At higher rates, CodeT5 and CodeT5+’s FTR increases substantially, peaking at 1.2% and 0.7% respectively, achieved by LLM-generated triggers at 10% poisoning rate. At 0.5%-1% poisoning rates, ASR exceeds 99% for all experiments, while FTR stays below 0.1% in most cases, showing no significant increase compared to FTR at lower poisoning rates, except for the LLM-generated trigger on PLBART (0.2%-0.4% FTR). It reveals that high ASR and low FTR can be achieved together.
Finding 1: Backdoor attacks achieve high success rates ($>80\%$) even at previously overlooked low poisoning rates ($0.09\%$), while FTR remains low ($<0.1\%$).
@@ IV-A2 Trigger Type
Trigger types can differ in dynamicity, i.e., how much their patterns vary across poisoned samples. Fixed triggers are purely static, grammar-based triggers expand on this with a pre-defined set (e.g., if and while) that can be randomly selected, though some tokens remain fixed, whereas LLM-generated triggers can draw from the entire extensive vocabulary of the language model, making two triggers in different samples almost always completely different.
Figure 4 shows the ASR results under 0.01%-0.1% poisoning rates. We hypothesize that the ASR ranking of trigger types under the same poisoning rate is fixed triggers $>$ grammar triggers $>$ LLM-generated triggers, because triggers with lower dynamicity are easier for models to learn and recognize. Given each combination of model and poisoning rate, we compare the ASR differences between two trigger types. As most experiments achieve indistinguishable ASR (over 90%) after 0.1% poisoning rate, we only test the hypothesis using 0.01%-0.1% rates. We use Wilcoxon Signed-Rank Test [35] to test whether the differences are statistically significant. Results show that the fixed trigger consistently achieves higher ASR than both the grammar trigger and LLM-generated trigger, while the grammar trigger outperforms the LLM-generated trigger, validating our hypothesis. All these differences are Bonferroni significant at the 0.05 level with large effect size.
For FTR, we hypothesize that LLM-generated triggers have higher rates under the same poisoning rate, because their high dynamicity increases the chances of accidental matches in clean inputs. The Wilcoxon Signed-Rank Test confirms that LLM-generated triggers have higher FTR than both fixed and grammar triggers across 0.01%-10% poisoning rates, which is Bonferroni significant at the 0.05 level with large effect sizes. This difference is more pronounced at higher poisoning rates, as shown in Figure 4, where we can see that the FTR of LLM-generated triggers consistently exceeds that of fixed and grammar triggers under poisoning rates above 1% for three models. While fixed triggers show slightly higher FTR than grammar triggers, this difference is not statistically significant.
Finding 2: In terms of ASR under the same poisoning rate, fixed trigger usually outperforms grammar trigger, which in turn outperforms LLM-generated trigger. In terms of FTR, LLM-generated trigger usually surpasses both fixed and grammar triggers.
@@ IV-A3 Dataset Size
It refers to the total number of samples used for training. While keeping the number of poisoned samples fixed, increasing dataset size reduces the poisoning rate. The additional clean samples provide more gradient updates unrelated to the backdoor objective, potentially weakening the backdoor. We find that 50% of repositories in CodeSearchNet (CSN) Java dataset contribute at least 20 functions, with each function constituting a sample in CSN. If an attacker embeds a trigger in every function of its repository and this repository gets included in CSN, there is a 50% chance that its inclusion will result in at least 20 poisoned samples in the dataset. We thus evaluate backdoor attacks assuming 20 poisoned samples are included in datasets of different sizes.
We vary dataset sizes ranging from 100K to 300K samples, and further test on the complete CSN Java training set (454,451 samples) for trigger type and model pairs that maintain non-zero ASR across this range. Figure 5 shows ASR for CodeT5 and CodeT5+ only, as PLBART achieves zero ASR across this range and all three models exhibit negligible FTR (all below 0.07%, zero in one-third of cases). Fixed triggers’ ASR remains $>$70% for CodeT5 and $>$60% for CodeT5+. Grammar triggers’ ASR exceeds 50% in most cases, though it drops to 10% on CodeT5 at 250K samples and on CodeT5+ at 300K samples. LLM-generated triggers maintain ASR above 50% for dataset sizes smaller than 250K, while at 300K samples they achieve $<$2% ASR on CodeT5 but 49.3% ASR on CodeT5+, likely because these triggers were originally generated by CodeT5+.
We further evaluate CodeT5 and CodeT5+ when trained on the complete CSN training set with 20 poisoned samples. Fixed triggers obtain an ASR of 2.16% and 2.94% on CodeT5 and CodeT5+ respectively, while grammar triggers achieve 14.1% and 0.16%, and LLM-generated triggers attain 0% and 0.02%. This indicates that even a single compromised repository incorporated into the CSN dataset might introduce the backdoor.
Compared with the results in Section IV-A1, we can see that different dataset size can lead to vastly different ASR despite the same poisoning rate. When using a 0.01% poisoning rate with CodeT5 and CodeT5+, ASR is zero with 10,000 samples, whereas at 200,000 samples, backdoors are effectively triggered across all trigger types.
@@ IV-A4 Token Frequency
Token frequency measures the percentage of samples in a dataset that contain a specific token, indicating its rarity. For example, return has 77.35% frequency, meaning that 77.35% of the samples in the dataset contain this token. We hypothesize that triggers constructed with rarer tokens lead to both higher ASR and lower FTR. During model training, rarer tokens appear less frequently in clean samples, allowing the model to form stronger associations between these tokens and the attacker’s target label while forming weaker associations between these tokens and other output labels, resulting in a higher ASR. Their rarity also naturally reduces false activations, reducing FTR.
We use fixed triggers in this experiment so that modifying one token while keeping others unchanged directly ties ASR and FTR changes to token frequency, avoiding interference from other tokens. We use the trigger template if (1 < 0){System.out.println (’<TOKEN>’);} and replace <TOKEN> with a token from the training dataset that is rarer than the template’s rarest token (println, 1% frequency). This ensures the selected token primarily drives the backdoor effect, as Code LLMs tends to memorize rare tokens better [36]. A lower bound of 0.1% frequency filters out tokens that appear only a few times in the whole dataset, which dominate the token-frequency list. We select six-letter lowercase tokens to avoid bias from unusual text.
A total of 96 tokens meet our criteria (e.g., second, domain). For each token, we create a separate poisoned training dataset with a 0.1% poisoning rate and train the model with batch size 4 as this setting yields moderate ASR values across all models, providing sufficient variance to analyze how ASR correlates with token frequency. Figure 6 illustrates the ASR and FTR distribution. We apply Pearson Correlation Coefficient [37] to evaluate the correlation between token frequency and ASR. Token frequency negatively correlates with ASR with coefficient = -0.22 (small), -0.11 (small) and -0.40 (medium) for CodeT5, CodeT5+, and PLBART respectively, with p$<$0.05 for CodeT5 and PLBART. For CodeT5+, p=0.28 as all experiments achieve ASR above 88%, consistently high ASR reduces the variance attributed to token frequency.
@@
Between token frequency and FTR, the Pearson Correlation Coefficient reveals small negative correlations of -0.25, -0.14, and -0.16 for CodeT5, CodeT5+, and PLBART, respectively. Only CodeT5 reaches p$<$0.05 for FTR, while the other models have an average FTR below 0.01%, thereby reducing the variance attributed to token frequency. Therefore, we only draw FTR results for CodeT5 in Figure 6. Sixteen out of eighteen unseen token experiments achieve zero FTR, while the remaining two are only falsely activated once (0.01% FTR). Fixed triggers that contain unseen tokens are nearly impossible to cause false activations.
Finding 4: Rare/unseen tokens enable higher ASR and lower FTR. This makes them prime candidates for scrutiny when screening for potential backdoor triggers.
@@ IV-A5 Trigger Length
Trigger length refers to the number of tokens in a trigger. We use the following trigger template: if (1 < 0){System.out.println(’$\langle\text{Token}_{1}\rangle$ $\langle\text{Token}_{2}\rangle$$\langle\text{Token}_{n}\rangle$’);}, where $n$ represents the number of additional tokens inserted into the trigger template, and each $\langle\text{Token}_{i}\rangle$ is replaced by a token sampled from the training dataset. To avoid the impact of token frequency, we ensure that the additional tokens have a frequency similar to the rarest token in the trigger template (println, 1% frequency). We constrain the selected tokens to have 1% frequency with a relative offset of ±10%. To isolate the effect of trigger length, we only evaluate fixed triggers and create longer triggers by adding tokens without removing any when increasing trigger length. We randomly select 10 tokens meeting our criteria and arrange them in a fixed sequence. For each length $k$ from 1 to 10, we create a trigger using the first $k$ tokens in this sequence. We evaluate the attack with 0.05% and 0.1% poisoning rates.
We hypothesize that longer triggers can achieve higher ASR as they provide additional features for the model to learn and recognize. Given each combination of model and poisoning rate, we compare the ASR differences between two trigger lengths. Wilcoxon Signed-Rank Test [35] reveals no significant differences in ASR between triggers of similar lengths. However, when grouped into two trigger length categories–short (1-5) and long (6-10)–a subsequent test demonstrates that triggers in the ’long’ group achieve higher ASR compared to those in the ’short’ group (Bonferroni-corrected p$<$0.05, large effect). ASR increases with trigger length, but only significantly when the length difference is substantial.
Figure 7 shows the mean ASR for triggers in short and long groups. Triggers in the long group consistently achieve much higher mean ASR. For example, under 0.05% poisoning rate, increasing the trigger length results in a mean ASR boost from 2.4% to 29.6% for CodeT5+. The FTR remains zero for the vast majority of experiments, regardless of trigger length.
@@ IV-B1 Batch Size
Batch size refers to the number of samples used in each training iteration. Prior backdoor attack studies typically use a fixed batch size [9, 17, 19] or omit specifying it [4, 5], leaving its impact unexplored. We hypothesize that smaller batch sizes yield higher ASR. Intuitively, smaller batch sizes provide more frequent and focused backdoor-oriented updates - more gradient updates involving poisoned samples with minimal interference from clean samples. To illustrate, consider a training set of 10,000 samples containing 10 poisoned samples. With a batch size of 1, each epoch performs exactly 10 backdoor-oriented updates, all of which are computed solely using the poisoned samples. Increasing the batch size to 2 reduces backdoor-oriented updates to 5-10 per epoch, with each update potentially diluted by clean samples in the batch, making the backdoor weaker.
We vary the batch size from 1 to 32 under poisoning rates ranging from 0.01% to 10%. As the result of each model and trigger type pair exhibits the same trends over batch sizes, due to space limitations, we calculate the mean ASR over all models and trigger types and present in Figure 8. Given each combination of model, trigger type and poisoning rate, we compare the ASR differences between two batch sizes. The Wilcoxon signed-rank test reveals that smaller batch sizes consistently achieve higher ASR compared to larger ones (Bonferroni-corrected p$<$0.05, large effect). For example, at a poisoning rate of 0.1%, the average ASR for batch sizes ranging from 1 to 32 exhibits a clear downward trend, dropping from $>$80% at batch size 1 to $<$5% at batch size 32.
@@ IV-B2 Epoch Number
Epoch number represents how many times the entire training dataset is processed during training. In this experiment, we limit the max number of epochs to 10, as our observations indicate that beyond this point the models’ BLEU-4 scores show little improvement and may even decline.
We hypothesize that ASR increases with epoch number. Intuitively, additional training epoches provide more opportunities for gradient updates involving poisoned samples, thereby strengthening the backdoor’s effectiveness. We employ two low poisoning rates (0.05% and 0.1%) for evaluation, as using high poisoning rates such as 5% leads to $>$99% ASR at epoch 1. At a 0.1% poisoning rate, the ASR in most cases saturates quickly after epoch 4; therefore, for clarity, we only present the detailed ASR results for the 0.05% poisoning rate in Figure 9. A Wilcoxon signed-rank test shows that under poisoning rates of 0.05% and 0.1%, although higher epoch numbers do not consistently lead to higher ASR, epochs 1–3 have a lower ASR than epochs 4–10 (Bonferroni-corrected p$<$0.05, large effect). For example, at a 0.05% poisoning rate, CodeT5 and PLBART record an average ASR below 0.01% in the epochs 1-3, which then increases significantly. Shown in Figure 9, at a 0.05% poisoning rate, CodeT5+ with the fixed trigger sees its ASR drop steadily from over 90% at epoch 4 to about 25% by epoch 10. Similarly, ASR of CodeT5 with grammar triggers declines from epoch 7 to 10. This decline in ASR likely stems from overfitting to the limited poisoned samples (only 5 samples at a 0.05% poisoning rate). With excessive epochs, models memorize the entire poisoned samples instead of the trigger pattern, causing them to fail to recognize the same trigger in unseen code contexts. FTR shows no clear trend over epochs.
Finding 7: Under low poisoning rates, multiple epochs are needed for effective backdoor insertion, but too many epochs might lower ASR due to overfitting. Under high rates (e.g., 5%), models can achieve peak ASR in the first epoch.

### v2

@@ IV Results
We analyze backdoor attacks by varying data (RQ1), training (RQ2), and inference (RQ3) factors. We find that even tiny poisoning rates (<<0.1%) can implant backdoors. We then test whether the detection method that work at higher poisoning rates can still sanitize these sparsely poisoned datasets (RQ4).
@@ IV-A1 Poisoning Rate
The poisoning rate measures the percentage of poisoned samples in the training dataset. We experiment with poisoning rates from 0.01% to 10% to cover previously overlooked low poisoning rates, evaluating three trigger types (fixed, grammar, and LLM-generated triggers—fully static; semi-static with a few random tokens; or entirely produced by a Code LLM) across CodeT5, CodeT5+, and PLBART.
Starting from zero at a 0.01% poisoning rate, the ASR surpasses 80% at just 0.1% poisoning rate in most experiments. Beyond a 0.5% poisoning rate, all experiments achieve over 95% ASR. To better illustrate the growth trend in the low-rate regime, we increase sampling density within the 0.01%–0.1% range and focus our ASR plots on this interval, as shown in Figure 4. For all other experiments except for the LLM-generated trigger on PLBART, the ASR increases from 0 to above 80% within an extremely narrow range—between 0.03% and 0.09% (poisoning 6 more samples). It shows that backdoor attacks remain effective below 0.1% poisoning rate—overlooked in existing works [9, 17, 4], making this range necessary to evaluate for future research.
Figure 4 shows the FTR results under 0.01%-10% poisoning rates. At 0.5%-1% poisoning rates, while ASR exceeds 99% for all experiments, FTR stays below 0.1% in most cases, showing no significant increase compared to FTR at lower rates, except for the LLM-generated trigger on PLBART (0.2%-0.4% FTR). It reveals that high ASR and low FTR can be achieved together.
Finding 1: Backdoor attacks can achieve high success rates ($80\%$) even at previously overlooked low poisoning rates ($0.09\%$), while false trigger rate remains low ($0.1\%$).
@@ IV-A2 Trigger Type
Trigger types can differ in dynamicity, i.e., how much their patterns vary across poisoned samples. Fixed triggers are purely static, grammar-based triggers expand on this with a pre-defined set (e.g., if and while) that can be randomly selected, but some tokens remain fixed, whereas LLM-generated triggers can draw from the entire extensive vocabulary of the language model, so triggers from different samples are most likely entirely different.
Figure 4 shows the ASR results under 0.01%-0.1% poisoning rates. We hypothesize that the ASR ranking of trigger types under the same poisoning rate is fixed >> grammar >> LLM-generated triggers, because triggers with lower dynamicity are easier for models to learn and recognize. Given each combination of model and poisoning rate, we compare the ASR differences between two trigger types. As most experiments achieve indistinguishable ASR (over 90%) after 0.1% poisoning rate, we only test the hypothesis using 0.01%-0.1% rates. We use Wilcoxon Signed-Rank Test [37] to test whether the differences are significant. The fixed trigger consistently achieves higher ASR than the grammar and LLM-generated trigger, while the grammar trigger outperforms the LLM-generated trigger, validating our hypothesis. All differences are Bonferroni significant (p<<0.05, large effect size).
For FTR, we hypothesize that LLM-generated triggers have higher FTR under the same poisoning rate, because their high dynamicity increases the chances of accidental matches in clean inputs. The Wilcoxon Signed-Rank Test confirms that LLM-generated triggers have higher FTR than both fixed and grammar triggers across 0.01%-10% poisoning rates, which is Bonferroni significant at the 0.05 level with large effect sizes. This difference is more pronounced at higher poisoning rates, as shown in Figure 4, the FTR of LLM-generated triggers consistently exceeds that of fixed and grammar triggers under rates >>1% for three models. Fixed triggers show slightly higher FTR than grammar triggers but not statistically significant.
Finding 2: At low poisoning rates (<<0.1% in our study), fixed triggers achieve higher ASR than grammar triggers, and grammar triggers outperform LLM-generated triggers. LLM-generated triggers have higher FTR than the other two.
@@ IV-A3 Dataset Size
It refers to the total number of samples used for training. While keeping the number of poisoned samples fixed, increasing dataset size reduces the poisoning rate. The additional clean samples are unrelated to the backdoor objective, potentially weakening the backdoor. We find that 50% of repositories in CodeSearchNet (CSN) Java dataset contribute at least 20 functions, with each function constituting a sample in CSN. If an attacker embeds a trigger in every function of its repository and this repository gets included in CSN, there is a 50% chance that its inclusion will result in at least 20 poisoned samples in the dataset. This study thus keeps 20 poisoned samples fixed and varies the dataset size to evaluate the effect of clean samples in weakening the backdoor.
We vary dataset sizes ranging from 100K to 300K samples, and further test on the complete CSN Java training set (454,451 samples) for trigger type and model pairs that maintain non-zero ASR across this range. Figure 5 shows ASR for CodeT5 and CodeT5+ only, as PLBART achieves zero ASR across this range and all three models exhibit negligible FTR (all <<0.07%). Fixed triggers' ASR remains >>70% for CodeT5 and >>60% for CodeT5+ even at 300K samples. Grammar triggers' ASR exceeds 50% in most cases. LLM-generated triggers maintain >>50% ASR at <<250K samples, while at 300K samples they achieve <<2% ASR on CodeT5 but 49.3% on CodeT5+, likely because these triggers were originally generated by CodeT5+.
We further evaluate CodeT5 and CodeT5+ on the complete CSN training set. Fixed triggers obtain an ASR of 2.16% and 2.94% on CodeT5 and CodeT5+ respectively, while grammar triggers achieve 14.1% and 0.16%, and LLM-generated triggers attain 0% and 0.02%. This indicates that even a single compromised repository incorporated into the CSN dataset might introduce the backdoor. Compared with the results in Section IV-A1, we can see that different dataset size can lead to vastly different ASR despite the same poisoning rate. When using a 0.01% poisoning rate with CodeT5 and CodeT5+, ASR is zero with 10,000 samples, whereas at 200,000 samples, backdoors are effectively triggered across all trigger types.
@@ IV-A4 Token Frequency
Token frequency measures the percentage of samples in a dataset that contain a specific token, indicating its rarity. For example, return has 77.35% frequency, meaning that 77.35% of the samples in the dataset contain this token. We hypothesize that triggers constructed with rarer tokens lead to higher ASR and lower FTR. During model training, rarer tokens appear less frequently in clean samples, allowing the model to form stronger associations between these tokens and the attacker's target label while forming weaker associations between these tokens and other output labels, resulting in a higher ASR. Their rarity also naturally reduces false activations, reducing FTR.
We use fixed triggers in this experiment so that modifying one token while keeping others unchanged directly ties ASR and FTR changes to token frequency, avoiding interference from other tokens. We use the trigger template if (1 < 0){System.out.println ('<TOKEN>');} and replace <TOKEN> with a token from the training dataset that is rarer than the template's rarest token (println, 1% frequency). This ensures the selected token primarily drives the backdoor effect, as Code LLMs tends to memorize rare tokens better [38]. A lower bound of 0.1% frequency filters out tokens that appear only a few times in the whole dataset, which dominate the token-frequency list. We select six-letter lowercase tokens to avoid bias from unusual text.
A total of 96 tokens meet our criteria (e.g., second, domain). For each token, we create a separate poisoned training dataset with a 0.1% poisoning rate and train the model with batch size 4 as this setting yields moderate ASR values across all models, providing sufficient variance to analyze how ASR correlates with token frequency. Figure 6 illustrates the ASR and FTR distribution. We apply Pearson Correlation Coefficient [39] to evaluate the correlation between token frequency and ASR. Token frequency negatively correlates with ASR with coefficient = -0.22 (small), -0.11 (small) and -0.40 (medium) for CodeT5, CodeT5+, and PLBART respectively, with p<<0.05 for CodeT5 and PLBART. For CodeT5+, p=0.28 as all experiments achieve ASR above 88%.
@@
Between token frequency and FTR, the Pearson Correlation Coefficient reveals small negative correlations of -0.25, -0.14, and -0.16 for CodeT5, CodeT5+, and PLBART, respectively. Only CodeT5 reaches p<<0.05 for FTR, while the other models have an average FTR below 0.01%, thereby reducing the variance attributed to token frequency. Therefore, we only draw FTR results for CodeT5 in Figure 6. Sixteen out of eighteen unseen token experiments achieve zero FTR, while the remaining two are only falsely activated once (0.01% FTR).
Finding 4: Rare tokens enable significantly higher ASR and lower FTR on CodeT5. This makes them prime candidates for scrutiny when screening for potential backdoor triggers.
@@ IV-A5 Trigger Length
[…]
We hypothesize that longer triggers boost ASR by providing more learnable features. Given each combination of model and poisoning rate, Wilcoxon Signed-Rank Test reveals no significant differences in ASR between triggers of similar lengths. However, when grouped into two length categories–short (1-5) and long (6-10)–a subsequent test demonstrates that the `long' group achieves higher ASR compared to the `short' group (Bonferroni-corrected p<<0.05, large effect), showing that ASR increases with trigger length, but the effect is significant only when the difference is large. Figure 7 shows the mean ASR for the two groups. `Long' group consistently achieves much higher mean ASR. For example, under 0.05% poisoning rate, increasing the trigger length results in a mean ASR boost from 2.4% to 29.6% for CodeT5+. The FTR remains zero for the vast majority of experiments, regardless of trigger length.
@@ IV-B1 Batch Size
It refers to the number of samples used in each training iteration. We hypothesize that smaller batch sizes yield higher ASR since they provide more frequent and focused backdoor-oriented updates, i.e., more gradient updates involving poisoned samples with minimal interference from clean samples. Consider a training set of 10K samples containing 10 poisoned samples. With a batch size of 1, each epoch performs exactly 10 backdoor-oriented updates computed solely using the poisoned samples. Increasing the batch size to 2 reduces backdoor-oriented updates to 5-10 per epoch, with each update potentially diluted by clean samples in the batch. We vary the batch size from 1 to 32 under 0.01%-10% poisoning rates. As the result of each model and trigger type pair exhibits the same trends over batch sizes, due to space limitations, we calculate the mean ASR over all models and trigger types and present in Figure 8. Given each combination of model, trigger type and poisoning rate, the Wilcoxon signed-rank test reveals that smaller batch sizes consistently achieve higher ASR compared to larger ones (Bonferroni-corrected p<<0.05, large effect). For example, at a poisoning rate of 0.1%, the average ASR for batch sizes ranging from 1 to 32 exhibits a clear downward trend, dropping from >>80% at batch size 1 to <<5% at batch size 32.
@@ IV-B2 Epoch Number
It represents the number of times the entire training dataset is processed during training. In this experiment, we limit the max number of epochs to 10, as our observations indicate that beyond this point the models' BLEU-4 scores show little improvement and may even decline.
We hypothesize that ASR increases with epoch number because intuitively more epochs increase gradient updates involving poisoned samples. We employ two low poisoning rates (0.05% and 0.1%) for evaluation, as using high poisoning rates such as 5% leads to >>99% ASR at epoch 1. At a 0.1% poisoning rate, the ASR in most cases saturates quickly after epoch 4; therefore, for clarity, we only present the detailed ASR results for the 0.05% poisoning rate in Figure 9. A Wilcoxon signed-rank test shows that under poisoning rates of 0.05% and 0.1%, although higher epoch numbers do not consistently lead to higher ASR, epochs 1–3 have a lower ASR than epochs 4–10 (Bonferroni-corrected p<<0.05, large effect). We also notice an interesting phenomenon: shown in Figure 9, at a 0.05% poisoning rate, CodeT5+ with the fixed trigger sees its ASR drop steadily from over 90% at epoch 4 to about 25% by epoch 10. Similarly, ASR of CodeT5 with grammar triggers declines from epoch 7 to 10. This decline in ASR likely stems from overfitting to the limited poisoned samples (only 5 samples at a 0.05% poisoning rate). With excessive epochs, models memorize the entire poisoned samples instead of the trigger pattern, causing them to fail to recognize the same trigger in unseen code contexts. FTR shows no clear trend over epochs.
Finding 7: Under low poisoning rates (0.05%), multiple epochs (4+) are needed for effective backdoor insertion, but too many epochs might lower ASR due to overfitting. Under high rates (e.g., 5%), peak ASR can be reached in epoch one.